OpenStudy (anonymous):

what does z equal in this equation 5/6 = z/12

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

2 times 6 is 12 what is 2 times 5?

OpenStudy (anonymous):

hi 10

OpenStudy (misty1212):

bingo!

OpenStudy (anonymous):

easily, 5 divided by 6 is the same as Z divided by 12

OpenStudy (anonymous):

divide 5 by 6 and multiply by 12

OpenStudy (anonymous):

thank you can you help whith some more

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

Por misty helped too

OpenStudy (anonymous):

what is 1/8 times 2/9

OpenStudy (misty1212):

\[\frac{1}{8}\times \frac{2}{9}=\frac{1\times 2}{8\times 9}=\frac{1\times 1}{4\times 9}\]

OpenStudy (misty1212):

multiply gives \[\frac{1}{36}\]

OpenStudy (anonymous):

such smart

OpenStudy (anonymous):

what is 5/18 + 1/6

OpenStudy (anonymous):

lol misty is like step back

OpenStudy (misty1212):

\[\frac{5}{18}+\frac{1}{3}=\frac{5}{18}+\frac{3}{18}\]add up in the numerator

OpenStudy (anonymous):

lemme do this

OpenStudy (misty1212):

mmm ok

OpenStudy (anonymous):

Nooo YOU are saying that, I suck at fractions